OpenStudy (anonymous):

Captain Stacey is going to load her sailboat with jars of honey. She has room for 11 jars. She has allotted for 58 pounds of honey on her boat. She can choose from 3-pound jars (the variable j) or 7-pound jars (the variable h). Which equations would be a constraint to Captain Stacey maximizing her honey load.

OpenStudy (anonymous):

a) j - h = 11 b) j = 11 + h c) 3j = 58 + 7h d) 3j = 58 - 7h